3. Find exact solutions on the interval [0,21). 6cos2x + 7 sin x - 8 = 0
Answer #1

Solution 6 Cosa +7 Simu - 8=0 we know that Sinats n =) Cas r=1-S1m² thus 6( 1- Sm?v) + 7 Simu -8=0 6-6 Smx + 7 Simo-8=0 - 6 S